OpenStudy (anonymous):

Ravi is making a mixture of sand, cement and water. He adds 6 kg of cement and 4 litres of water to every 10 kg of sand. Which of these statements is not correct? 1) The water is 20 % of the mixture. 2) The ratio of sand to cement to water is 5:2:3. 3) To make 30 kg of mixture, Ravi needs 9 kg of cement. 4) If he uses 4.5 kg of cement, then he needs 3 L of water.

OpenStudy (whpalmer4):

1 l of water is 1 kg 6kg cement + 4kg water + 10kg of sand makes 20kg of mixture 4kg water/20kg mixture = 1/5 = 0.2 = 20% 10 kg sand : 6 kg cement : 4 kg water : factor out common factor of 2 gives 5:3:2 to make 30 kg of mixture, he need 6 kg/20*30 = 9 kg of cement if he uses 4.5 kg of cement, he needs 4.5*2/3 = 3 kg = 3l of water So the ratio of sand to cement to water answer is the incorrect statement.